Can you help me on this inverse funtions question:
1)Determine the inverse of each of the following functions. State the domain and range of the inverse. For which of these is the inverse also a function?
a)f(x)=2(x-2)^2

Answer by stanbon(75887)

Determine the inverse of each of the following functions. State the domain and range of the inverse. For which of these is the inverse also a function?
a)f(x)=2(x-2)^2
-------------------
1st: Interchange x and y to get:
x = 2(y-2)^2
----
2nd: Solve for "y":
(y-2)^2 = (x/2)
y-2 = sqrt(x/2) or y-2 = -sqrt(x/2)
y = 2+sqrt(x/2) or y = 2-sqrt(x/2)
---------------
The inverse is not a function because there are two y-values for
each value of "x".
===============================
Domain of the inverse: Since x/2 must not be negative the domain is
All Real Numbers greater than or equal to zero.
---
Range of the inverse: Since the domain of the given function is
all Real Numbers, the range of the inverse is all Real Numbers.
==================================================================
Cheers,
Stan H.
========================